OpenStudy (anonymous):

solve the system by the addition method : -3x+7y=-9 27x-63y=82

OpenStudy (callisto):

I doubt if there is a possible solution -3x+7y=-9 -(1) 27x-63y=82 -(2) (1)x9 +(2) 9(-3x+7y) +(27x-63y) = -9 (9) +82 -27x +63y +27x -63y = -81 +82 0 = -1 Alright that's what i got :S

OpenStudy (anonymous):

what does x = and what does y =

OpenStudy (callisto):

-3x+7y=-9 => x = (9+7y) /3 -(1) 27x-63y=82 -(2) Sub. (1) into (2) 27 [(9+7y) /3] - 63y =82 81 +63y - 63y =82 81 = 82 Alright what do you think for the answer then?

OpenStudy (anonymous):

okay thanks

OpenStudy (callisto):

Welcome